# NeuralSurv: Deep Survival Analysis with Bayesian Uncertainty Quantification

### v1

@@ Appendix A Review of Survival Analysis
This appendix offers a concise summary of the survival-analysis framework on which our approach is built. For an in-depth review, the reader is referred to [27].
@@
To handle censoring uniformly, we introduce a censoring time $C$ and record the observed time $y=\min(T,\,C)$. The event indicator can then be written succinctly as $\delta=\mathbbm{1}_{\{T\leq C\}}$. Throughout, we assume noninformative right-censoring, i.e. conditional on the covariates, the censoring time is independent of the event time: $C\perp T\mid\mathbf{x}$. Conditional on $\mathbf{x}$, we let the event time $T$ have cumulative distribution function $F(t\mid\mathbf{x})$ and probability density function $f(t|\mathbf{x})$ such that
@@ Appendix B Review of Pólya-Gamma Random Variables
We follow [34] in defining the family of Pólya–Gamma distributions and their properties.
@@
[…]
@@
[…]
@@
The Pólya–Gamma density (19) can be re-written in the form
$p_{\text{PG}}(\omega\mid b,c)=\exp\left(-\frac{c^{2}}{2}\omega\right)\,(\cosh(c/2))^{b}\,p_{\text{PG}}(\omega\mid b,0).$ (20)
@@
[…]
In particular, the first moment is obtained by differentiating (21) with respect to $\xi$ at $\xi=0$:
$\mathbb{E}_{\omega\sim p_{\text{PG}}(\omega\mid b,c)}[\omega]=\frac{b}{2c}\tanh\left(\frac{c}{2}\right).$ (22)
@@
[…]
@@
[…]
@@ Appendix C Review of Poisson processes
This appendix briefly summarizes the properties of a Poisson process that are most relevant to our analysis. For a more comprehensive treatment, see Chapters 3 and 5 of [22].
@@
Given a point process $\Psi$, we denote its path measure — that is, the probability measure induced on its sample-path space — by $\mathbb{P}_{\Psi}$. If the intensity function $\lambda(z)$ is constant, $\lambda(z)\equiv\lambda$, then $\Psi$ is called homogeneous; otherwise, it is inhomogeneous. We now extend the concept of a Poisson process by incorporating additional random attributes, known as marks.
@@
Next, we present Campbell’s Theorem, which describes the law of sums taken over the points of a Poisson process (see [22, Sec. 3.2]).
@@
Appendix D Obtaining the normalizing function $Z(t,\mathbf{x})$
In this appendix we derive an efficient approximation for the normalizing constant
$Z(t,\mathbf{x})=\mathbb{E}_{\boldsymbol{\theta}\sim p_{\boldsymbol{\theta}}}\left[\sigma(g^{\text{lin}}(t,\mathbf{x};\boldsymbol{\theta}))\right],$ (24)
@@ Appendix D Obtaining the normalizing function $Z(t,\mathbf{x})$
$\mathbb{E}_{X\sim\mathcal{N}(\mu,\sigma^{2})}[\sigma(X)]\approx\sigma\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right).$ (25)
We can apply the result in (25) to the normal random variable $g^{\text{lin}}(t,\mathbf{x};\boldsymbol{\theta})$ and approximate $Z(t,\mathbf{x})$ as
@@
Since in (24) we are taking the expectation under the prior $p_{\boldsymbol{\theta}}(\boldsymbol{\theta})$, it is natural to linearize around the prior mean, therefore, we set $\boldsymbol{\theta}^{\star}=\boldsymbol{0}$.
@@ Appendix E Combining Variational Inference with Poisson Processes
In this appendix, we outline how our variational-inference framework integrates marked Poisson processes — an essential part in the mean-field variational approximation of Section 4.1. For a fully rigorous, measure-theoretic treatment, the reader is referred to Brémaud’s text [5]. Our development relies in particular on Theorem T10 in Chapter VIII of that book, which shows how the law of a marked Poisson process arises via a change of measure using the appropriate Radon–Nikodym derivative.
@@
Let $\boldsymbol{\Psi}=(\Psi_{1},\dots,\Psi_{N})$ be $N$ independent marked Poisson processes, where each $\Psi_{i}$ is defined on the product space $[0,y_{i}]\times\mathbb{R}_{+}$. We define $\mathbb{P}_{\boldsymbol{\Psi},*}$ to be their joint law where each $\Psi_{i}$ has intensity
$\lambda_{*,i}(t,\omega)\;=\;t^{\rho-1}p_{\mathrm{PG}}(\omega\mid 1,0)\quad\text{for all }(t,\omega)\in[0,y_{i}]\times\mathbb{R}_{+}.$ (26)
Next, let $\gamma_{i}^{\mathbb{Q}}(t)$ be a deterministic function on $[0,y_{i}]$ and let $h_{i}^{\mathbb{Q}}(t,\omega)$ be a deterministic density on $[0,y_{i}]\times\mathbb{R}_{+}$ satisfying
$\int_{0}^{\infty}h_{i}^{\mathbb{Q}}(t,\omega)p_{\text{PG}}(\omega|1,0)\mathrm{d}\omega=1\quad\quad\text{and}\quad\int_{0}^{y_{i}}\gamma_{i}^{\mathbb{Q}}(t)t^{\rho-1}\mathrm{d}t<\infty$ (27)
@@
By Theorem T10.b [5, Chapter VIII], whenever $\mathbb{E}_{\boldsymbol{\Psi}\sim\mathbb{P}_{\boldsymbol{\Psi},*}}[L(\boldsymbol{\Psi})]=1$, the measure $\mathbb{Q}_{\boldsymbol{\Psi}}(\boldsymbol{\Psi})$ defined by $\frac{\mathrm{d}\mathbb{Q}_{\boldsymbol{\Psi}}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi})=L(\boldsymbol{\Psi})$ is exactly the law under which each $\Psi_{i}$ is a marked Poisson process on $[0,y_{i}]\times\mathbb{R}_{+}$ with intensity $\lambda^{\mathbb{Q}}_{i}(t,\omega)$. The above result underpins the analysis in Appendix G.2, where we show that the optimal variational measure $\mathbb{Q}_{\boldsymbol{\Psi}}$ coincides with the law of a collection of independent marked Poisson processes.
@@
$\frac{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi}|\phi}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi})=\prod_{i=1}^{N}\left(\prod_{(t,\omega)_{j}\in\Psi_{i}}\frac{\phi}{Z(t_{j},\mathbf{x}_{i})}\right)\exp\left(\int_{0}^{y_{i}}\int_{0}^{\infty}\left(\lambda_{*,i}(t,\omega)-\lambda_{i}(t,\omega;\phi)\right)\mathrm{d}\omega\mathrm{d}t\right).$ (28)
@@ F.1 Approximating the Log Likelihood distribution
Our aim is to approximate the log likelihood density $\log p(\mathcal{D}\mid\mathbf{X},g(\cdot;\boldsymbol{\theta}),\phi)$. In order to do so, we introduce a variational distribution ˘$\breve{\mathbb{Q}}(\boldsymbol{\omega},\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi)$ to approximate the true distribution $\mathbb{P}(\boldsymbol{\omega},\boldsymbol{\Psi}\mid\mathcal{D},\mathbf{X},g(\cdot;\boldsymbol{\theta}),\phi)$. Such variational distribution differs from the one used for full-model inference in Section 4.3 because it is conditioned on the values of $\boldsymbol{\theta}$ and $\phi$. Hence, we adopt the notation $\breve{\mathbb{Q}}$ (instead of $\mathbb{Q}$) to highlight this difference. We restrict our search to distributions that satisfy the following mean-field factorization:
˘$\breve{\mathbb{Q}}(\boldsymbol{\omega},\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi)=\breve{\mathbb{Q}}_{\boldsymbol{\omega}\mid\boldsymbol{\theta},\phi}(\boldsymbol{\omega}\mid\boldsymbol{\theta},\phi)\times\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi}(\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi).$
Here, we take ˘$\breve{\mathbb{Q}}_{\boldsymbol{\omega}\mid\boldsymbol{\theta},\phi}(\boldsymbol{\omega}\mid\boldsymbol{\theta},\phi)$ to admit the density $\breve{q}_{\boldsymbol{\omega}\mid\boldsymbol{\theta},\phi}(\boldsymbol{\omega}\mid\boldsymbol{\theta},\phi)$ with respect to the Lebesgue measure $\mathrm{d}\boldsymbol{\omega}$.
For the marked point process component, we assume that the variational law ˘$\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi}$ is absolutely continuous with respect to $\mathbb{P}_{\boldsymbol{\Psi},*}$, so that it admits a strictly positive Radon–Nikodym derivative $\frac{\mathrm{d}\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}$ which satisfies the normalization $\mathbb{E}_{\boldsymbol{\Psi}\sim\mathbb{P}_{\boldsymbol{\Psi},*}}\left[\frac{\mathrm{d}\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi})\right]=1$. These two conditions guarantee that ˘$\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi}$ is indeed a probability measure on the space of marked point-process paths.
@@
$\log p(\mathcal{D}\mid\mathbf{X},g(\cdot;\boldsymbol{\theta}),\phi)\ =\\ D_{\text{KL}}\Big{(}\breve{\mathbb{Q}}_{\boldsymbol{\omega},\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi}(\boldsymbol{\omega},\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi)\>||\>\mathbb{P}(\boldsymbol{\omega},\boldsymbol{\Psi}\mid\mathcal{D},\mathbf{X},g(\cdot;\boldsymbol{\theta}),\phi)\Big{)}+\breve{\mathcal{L}}_{\text{ELBO}},$ (29)
@@
˘$\breve{\mathcal{L}}_{\text{ELBO}}:=\mathbb{E}_{\boldsymbol{\omega}\sim\breve{q}_{\boldsymbol{\omega}\mid\boldsymbol{\theta},\phi},\boldsymbol{\Psi}\sim\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi}}\left[\log\frac{p(\mathcal{D}\mid\mathbf{X},g(\cdot;\boldsymbol{\theta}),\phi,\boldsymbol{\omega},\boldsymbol{\Psi})\>p_{\boldsymbol{\omega}}(\boldsymbol{\omega})\>\frac{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi}\mid\phi}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi})}{\breve{q}_{\boldsymbol{\omega}\mid\boldsymbol{\theta},\phi}(\boldsymbol{\omega}\mid\boldsymbol{\theta},\phi)\>\frac{\mathrm{d}\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi)}\right],$ (30)
and where $\frac{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi}\mid\phi}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}$ is the Radon-Nykodim derivative of the true conditional law $\mathbb{P}_{\boldsymbol{\Psi}\mid\phi}$ with respect to $\mathbb{P}_{\boldsymbol{\Psi},*}$, cf.  (28).
@@
When the variational distribution ˘$\breve{\mathbb{Q}}(\boldsymbol{\omega},\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi)$ matches the true posterior $\mathbb{P}(\boldsymbol{\omega},\boldsymbol{\Psi}\mid\mathcal{D},\mathbf{X},g(\cdot;\boldsymbol{\theta}),\phi)$, the KL divergence term in (29) vanishes. Consequently, the ELBO becomes equal to the marginal log-likelihood, and maximizing the ELBO is equivalent to maximizing log-likelihood directly. In practice, we minimize the KL divergence so that our ELBO provides the closest possible lower bound to the true log-likelihood. Therefore, in order to obtain the closest lower bound to to the log-likelihood $\log p(\mathcal{D}\mid\mathbf{X},g(\cdot;\boldsymbol{\theta}),\phi)$ we must find the distribution ˘$\breve{\mathbb{Q}}(\boldsymbol{\omega},\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi)$ which minimizes the KL divergence in (29).
Using standard mean-field variational inference techniques (see, e.g., Chapter 10.1 of [3]), the optimal distribution for the latent variables $\boldsymbol{\omega}$ given $(\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)})$ is obtained by computing the expectation of the joint log-density with respect to the other variational factors, that is
@@
$\>\>\mathbb{E}_{\boldsymbol{\Psi}\sim\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)}}}\left[\log p(\mathcal{D}\mid\mathbf{X},g(\cdot;\boldsymbol{\theta}^{(\ell)}),\phi^{(\ell)},\boldsymbol{\omega},\boldsymbol{\Psi})+\log p_{\boldsymbol{\omega}}(\boldsymbol{\omega})+\log\frac{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi}\mid\phi^{(\ell)}}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi})\right]+\text{const.}$.
@@
$\>\>\mathbb{E}_{\boldsymbol{\omega}\sim\breve{q}_{\boldsymbol{\omega}\mid\boldsymbol{\theta},\phi}}\left[\log p(\mathcal{D}\mid\mathbf{X},g(\cdot;\boldsymbol{\theta}^{(\ell)}),\phi^{(\ell)},\boldsymbol{\omega},\boldsymbol{\Psi})+\log p_{\boldsymbol{\omega}}(\boldsymbol{\omega})+\log\frac{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi}\mid\phi^{(\ell)}}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi})\right]+\text{const.}$.
@@
$\breve{q}_{\boldsymbol{\omega}\mid\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)}}(\boldsymbol{\omega}\mid\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)})=\prod_{i=1}^{N}\breve{q}_{\omega_{i}\mid\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)}}(\omega_{i}\mid\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)})=\prod_{i=1}^{N}p_{\text{PG}}\left(\omega_{i}\mid 1,\breve{c}^{(\ell)}_{i}\right),$ (31)
@@
$\breve{c}^{(\ell)}_{i}=\delta_{i}\>|g(y_{i},\mathbf{x}_{i};\boldsymbol{\theta}^{(\ell)})|.$ (32)
By mirroring the derivation in Appendix G.2, one shows that the optimal measure ˘$\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi}(\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi)$ is exactly the law under which each $\Psi_{i}$, for $i=1,\ldots,N$, is a marked Poisson process on $[0,y_{i}]\times\mathbb{R}_{+}$ with intensity

### v2

@@ Appendix A Review of Survival Analysis
This appendix offers a concise summary of the survival-analysis framework on which our approach is built. For an in-depth review, the reader is referred to [Lawless_survival].
@@
To handle censoring uniformly, we introduce a censoring time $C$ and record the observed time $y=\min(T,\,C)$. The event indicator can then be written succinctly as $\delta=\mathbbm{1}_{\{T\leq C\}}$. Throughout, we assume noninformative right‑censoring, i.e. conditional on the covariates, the censoring time is independent of the event time: $C\perp T\mid\mathbf{x}$. Conditional on $\mathbf{x}$, we let the event time $T$ have cumulative distribution function $F(t\mid\mathbf{x})$ and probability density function $f(t|\mathbf{x})$ such that
@@ Appendix B Review of Pólya-Gamma Random Variables
We follow [Polson01122013] in defining the family of Pólya–Gamma distributions and their properties.
@@
$\omega\stackrel{{\scriptstyle d}}{{=}}\frac{1}{2\pi^{2}}\sum_{k=1}^{\infty}\frac{g_{k}}{\left(k-\frac{1}{2}\right)^{2}+\frac{c^{2}}{4\pi^{2}}},\quad\text{with }g_{k}\overset{\text{i.i.d.}}{\sim}\text{Gamma}(b,1).$ (A1)
@@
$\left[\cosh\left(\frac{c}{2}\right)\right]^{-b}=\int_{0}^{\infty}\exp\left(-\frac{c^{2}}{2}\omega\right)\,p_{\text{PG}}(\omega\mid b,0)\,d\omega.$
@@
The Pólya–Gamma density (A1) can be re-written in the form
$p_{\text{PG}}(\omega\mid b,c)=\exp\left(-\frac{c^{2}}{2}\omega\right)\,(\cosh(c/2))^{b}\,p_{\text{PG}}(\omega\mid b,0).$ (A2)
@@
$\int_{0}^{\infty}e^{\xi\omega}\,p_{\text{PG}}(\omega\mid b,c)\,d\omega=\frac{\cosh^{b}(c/2)}{\cosh^{b}\!\left(\frac{1}{2}\sqrt{c^{2}-2\xi}\,\right)}.$ (A3)
In particular, the first moment is obtained by differentiating (A3) with respect to $\xi$ at $\xi=0$:
$\mathbb{E}_{\omega\sim p_{\text{PG}}(\omega\mid b,c)}[\omega]=\frac{b}{2c}\tanh\left(\frac{c}{2}\right).$ (A4)
@@
$\frac{e^{\psi a}}{(1+e^{\psi})^{b}}=2^{-b}e^{\kappa\psi}\int_{0}^{\infty}\exp\left({-\frac{\omega\psi^{2}}{2}}\right)\,p_{\text{PG}}(\omega\mid b,0)\,d\omega,$
@@
$\sigma(z)=\frac{e^{\frac{z}{2}}}{2\cosh(\frac{z}{2})}=\int^{\infty}_{0}e^{f(\omega,z)}p_{\text{PG}}(\omega\mid 1,0)d\omega.$ (A5)
@@ Appendix C Review of Poisson Processes
This appendix briefly summarizes the properties of a Poisson process that are most relevant to our analysis. For a more comprehensive treatment, see Chapters 3 and 5 of [Kingman1992].
@@
Given a point process $\Psi$, we denote its path measure — that is, the probability measure induced on its sample‑path space — by $\mathbb{P}_{\Psi}$. If the intensity function $\lambda(z)$ is constant, $\lambda(z)\equiv\lambda$, then $\Psi$ is called homogeneous; otherwise, it is inhomogeneous. We now extend the concept of a Poisson process by incorporating additional random attributes, known as marks.
@@
Next, we present Campbell’s Theorem, which describes the law of sums taken over the points of a Poisson process (see [Kingman1992, Sec. 3.2]).
@@
Appendix D Obtaining the Normalization Factor $Z(t,\mathbf{x})$
In this appendix we derive an efficient approximation for the normalization factor
$Z(t,\mathbf{x})=\mathbb{E}_{\boldsymbol{\theta}\sim p_{\boldsymbol{\theta}}}\left[\sigma(g^{\text{lin}}(t,\mathbf{x};\boldsymbol{\theta}))\right],$ (A6)
@@ Appendix D Obtaining the Normalization Factor $Z(t,\mathbf{x})$
$\mathbb{E}_{X\sim\mathcal{N}(\mu,\sigma^{2})}[\sigma(X)]\approx\sigma\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right).$ (A7)
We can apply the result in (A7) to the normal random variable $g^{\text{lin}}(t,\mathbf{x};\boldsymbol{\theta})$ and approximate $Z(t,\mathbf{x})$ as
@@
Since in (A6) we are taking the expectation under the prior $p_{\boldsymbol{\theta}}(\boldsymbol{\theta})$, it is natural to linearize around the prior mean, therefore, we set $\boldsymbol{\theta}^{\star}=\boldsymbol{0}$.
@@ Appendix E Combining Variational Inference with Poisson Processes
In this appendix, we outline how our variational‐inference framework integrates marked Poisson processes — an essential part in the mean‐field variational approximation of Section 4.1. For a fully rigorous, measure‐theoretic treatment, the reader is referred to Brémaud’s text [Bremaud-Point-Process]. Our development relies in particular on Theorem T10 in Chapter VIII of that book, which shows how the law of a marked Poisson process arises via a change of measure using the appropriate Radon–Nikodym derivative.
@@
Let $\boldsymbol{\Psi}=(\Psi_{1},\dots,\Psi_{N})$ be $N$ independent marked Poisson processes, where each $\Psi_{i}$ is defined on the product space $0,y_{i}]\times\mathbb{R}_{+}$. We define $\mathbb{P}_{\boldsymbol{\Psi},*}$ to be their joint law where each $\Psi_{i}$ has intensity
$\lambda_{*,i}(t,\omega)\;=\;t^{\rho-1}p_{\mathrm{PG}}(\omega\mid 1,0)\quad\text{for all }(t,\omega)\in[0,y_{i}]\times\mathbb{R}_{+}.$ (A8)
Next, let $\gamma_{i}^{\mathbb{Q}}(t)$ be a deterministic function on $0,y_{i}]$ and let $h_{i}^{\mathbb{Q}}(t,\omega)$ be a deterministic density on $0,y_{i}]\times\mathbb{R}_{+}$ satisfying
$\int_{0}^{\infty}h_{i}^{\mathbb{Q}}(t,\omega)p_{\text{PG}}(\omega|1,0)\mathrm{d}\omega=1\quad\quad\text{and}\quad\int_{0}^{y_{i}}\gamma_{i}^{\mathbb{Q}}(t)t^{\rho-1}\mathrm{d}t<\infty$ (A9)
@@
By Theorem T10.b [Bremaud-Point-Process, Chapter VIII], whenever $\mathbb{E}_{\boldsymbol{\Psi}\sim\mathbb{P}_{\boldsymbol{\Psi},*}}[L(\boldsymbol{\Psi})]=1$, the measure $\mathbb{Q}_{\boldsymbol{\Psi}}(\boldsymbol{\Psi})$ defined by $\frac{\mathrm{d}\mathbb{Q}_{\boldsymbol{\Psi}}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi})=L(\boldsymbol{\Psi})$ is exactly the law under which each $\Psi_{i}$ is a marked Poisson process on $0,y_{i}]\times\mathbb{R}_{+}$ with intensity $\lambda^{\mathbb{Q}}_{i}(t,\omega)$. The above result underpins the analysis in Appendix G.2, where we show that the optimal variational measure $\mathbb{Q}_{\boldsymbol{\Psi}}$ coincides with the law of a collection of independent marked Poisson processes.
@@
$\frac{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi}|\phi}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi})=\prod_{i=1}^{N}\left(\prod_{(t,\omega)_{j}\in\Psi_{i}}\frac{\phi}{Z(t_{j},\mathbf{x}_{i})}\right)\exp\left(\int_{0}^{y_{i}}\int_{0}^{\infty}\left(\lambda_{*,i}(t,\omega)-\lambda_{i}(t,\omega;\phi)\right)\mathrm{d}\omega\mathrm{d}t\right).$ (A10)
@@ F.1 Approximating the Log Likelihood distribution
[…]
$\breve{\mathbb{Q}}(\boldsymbol{\omega},\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi)=\breve{\mathbb{Q}}_{\boldsymbol{\omega}\mid\boldsymbol{\theta},\phi}(\boldsymbol{\omega}\mid\boldsymbol{\theta},\phi)\times\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi}(\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi).$
Here, we take $\breve{\mathbb{Q}}_{\boldsymbol{\omega}\mid\boldsymbol{\theta},\phi}(\boldsymbol{\omega}\mid\boldsymbol{\theta},\phi)$ to admit the density $\breve{q}_{\boldsymbol{\omega}\mid\boldsymbol{\theta},\phi}(\boldsymbol{\omega}\mid\boldsymbol{\theta},\phi)$ with respect to the Lebesgue measure $\mathrm{d}\boldsymbol{\omega}$.
For the marked point process component, we assume that the variational law $\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi}$ is absolutely continuous with respect to $\mathbb{P}_{\boldsymbol{\Psi},*}$, so that it admits a strictly positive Radon–Nikodym derivative $\frac{\mathrm{d}\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}$ which satisfies the normalization $\mathbb{E}_{\boldsymbol{\Psi}\sim\mathbb{P}_{\boldsymbol{\Psi},*}}\left[\frac{\mathrm{d}\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi})\right]=1$. These two conditions guarantee that $\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi}$ is indeed a probability measure on the space of marked point-process paths.
@@
$\log p(\mathcal{D}\mid\mathbf{X},g(\cdot;\boldsymbol{\theta}),\phi)\ =\\ D_{\text{KL}}\Big(\breve{\mathbb{Q}}_{\boldsymbol{\omega},\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi}(\boldsymbol{\omega},\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi)\>||\>\mathbb{P}(\boldsymbol{\omega},\boldsymbol{\Psi}\mid\mathcal{D},\mathbf{X},g(\cdot;\boldsymbol{\theta}),\phi)\Big)+\breve{\mathcal{L}}_{\text{ELBO}},$ (A11)
@@
$\breve{\mathcal{L}}_{\text{ELBO}}:=\mathbb{E}_{\boldsymbol{\omega}\sim\breve{q}_{\boldsymbol{\omega}\mid\boldsymbol{\theta},\phi},\boldsymbol{\Psi}\sim\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi}}\left[\log\frac{p(\mathcal{D}\mid\mathbf{X},g(\cdot;\boldsymbol{\theta}),\phi,\boldsymbol{\omega},\boldsymbol{\Psi})\>p_{\boldsymbol{\omega}}(\boldsymbol{\omega})\>\frac{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi}\mid\phi}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi})}{\breve{q}_{\boldsymbol{\omega}\mid\boldsymbol{\theta},\phi}(\boldsymbol{\omega}\mid\boldsymbol{\theta},\phi)\>\frac{\mathrm{d}\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi)}\right],$ (A12)
and where $\frac{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi}\mid\phi}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}$ is the Radon-Nykodim derivative of the true conditional law $\mathbb{P}_{\boldsymbol{\Psi}\mid\phi}$ with respect to $\mathbb{P}_{\boldsymbol{\Psi},*}$, cf.  (A10).
@@
When the variational distribution $\breve{\mathbb{Q}}(\boldsymbol{\omega},\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi)$ matches the true posterior $\mathbb{P}(\boldsymbol{\omega},\boldsymbol{\Psi}\mid\mathcal{D},\mathbf{X},g(\cdot;\boldsymbol{\theta}),\phi)$, the KL divergence term in (A11) vanishes. Consequently, the ELBO becomes equal to the marginal log-likelihood, and maximizing the ELBO is equivalent to maximizing log-likelihood directly. In practice, we minimize the KL divergence so that our ELBO provides the closest possible lower bound to the true log-likelihood. Therefore, in order to obtain the closest lower bound to to the log-likelihood $\log p(\mathcal{D}\mid\mathbf{X},g(\cdot;\boldsymbol{\theta}),\phi)$ we must find the distribution $\breve{\mathbb{Q}}(\boldsymbol{\omega},\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi)$ which minimizes the KL divergence in (A11).
Using standard mean-field variational inference techniques (see, e.g., Chapter 10.1 of [bishop_pattern_recognition]), the optimal distribution for the latent variables $\boldsymbol{\omega}$ given $(\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)})$ is obtained by computing the expectation of the joint log-density with respect to the other variational factors, that is
@@
$\>\>\mathbb{E}_{\boldsymbol{\Psi}\sim\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)}}}\left[\log p(\mathcal{D}\mid\mathbf{X},g(\cdot;\boldsymbol{\theta}^{(\ell)}),\phi^{(\ell)},\boldsymbol{\omega},\boldsymbol{\Psi})+\log p_{\boldsymbol{\omega}}(\boldsymbol{\omega})+\log\frac{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi}\mid\phi^{(\ell)}}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi})\right]+\text{const.}$
@@
$\>\>\mathbb{E}_{\boldsymbol{\omega}\sim\breve{q}_{\boldsymbol{\omega}\mid\boldsymbol{\theta},\phi}}\left[\log p(\mathcal{D}\mid\mathbf{X},g(\cdot;\boldsymbol{\theta}^{(\ell)}),\phi^{(\ell)},\boldsymbol{\omega},\boldsymbol{\Psi})+\log p_{\boldsymbol{\omega}}(\boldsymbol{\omega})+\log\frac{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi}\mid\phi^{(\ell)}}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi})\right]+\text{const.}$
@@
[…]
@@
$\breve{c}^{(\ell)}_{i}=\delta_{i}\>|g(y_{i},\mathbf{x}_{i};\boldsymbol{\theta}^{(\ell)})|.$ (A14)
By mirroring the derivation in Appendix G.2, one shows that the optimal measure $\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi}(\boldsymbol{\Psi}\mid\boldsymbol{\theta},\phi)$ is exactly the law under which each $\Psi_{i}$, for $i=1,\ldots,N$, is a marked Poisson process on $0,y_{i}]\times\mathbb{R}_{+}$ with intensity